OpenStudy (anonymous):

find the range of the following functions if the domain is {-2, -1, 0, 1, 2} f(x)=x(to the second power) + 1

jimthompson5910 (jim_thompson5910):

To find the range, simply plug in the given values of x to find corresponding values of f(x) So when x = -2, f(x) = x^2 + 1 f(-2) = (-2)^2 + 1 f(-2) = 4 + 1 f(-2) = 5 So when x = -2, f(x) = 5 So the same with the remaining x values to find corresponding f(x) values. These corresponding f(x) values will form the range.

OpenStudy (anonymous):

thank you!!!